# Qizhe Zhang1,2  Aosong Cheng1  Ming Lu3  Zhiyong Zhuo1  Minqi Wang1 Jiajun Cao1  Shaobo Guo2  Qi She2  Shanghang Zhang1

### v1

@@ Appendix D Efficiency Analysis with FlashAttention
FasterVLM (sdpa) 6.08 216.00 14.51 101.95 85.92
FasterVLM (flash attention) 101.35 85.87
@@
FasterVLM (sdpa) 4.16 144.00 14.46 93.57 86.20
FasterVLM (flash attention) 92.26 86.16
@@
FasterVLM (sdpa) 2.26 72.00 14.44 85.06 83.46
FasterVLM (flash attention) 84.03 83.42
@@
FasterVLM (sdpa) 1.13 29.00 14.54 79.11 75.85
FasterVLM (flash attention) 77.44 75.82
@@
FasterVLM (sdpa) 0.76 14.50 14.54 78.09 67.24
FasterVLM (flash attention) 77.15 67.22
Table 12: Efficiency comparison between FastV and FasterVLM under LLaVA-1.5-13B.
@@
FasterVLM (sdpa) 11.50 337.50 26.58 138.36 86.73
FasterVLM (flash attention) 137.95 86.72
@@
FasterVLM (sdpa) 7.76 225.00 26.31 124.13 86.05
FasterVLM (flash attention) 123.33 86.04
@@
FasterVLM (sdpa) 4.05 112.50 26.29 104.81 83.10
FasterVLM (flash attention) 103.57 83.09
@@
FasterVLM (sdpa) 1.86 45.31 26.27 94.80 74.71
FasterVLM (flash attention) 94.04 74.66
@@
FasterVLM (sdpa) 1.12 22.66 26.43 94.48 65.90
FasterVLM (flash attention) 93.73 65.79
Table 13: Efficiency comparison between FastV and FasterVLM under LLaVA-NeXT-7B.
@@
FasterVLM (sdpa) 32.35 1080.00 16.33 246.38 86.69
FasterVLM (flash attention) 232.07 86.66
@@
FasterVLM (sdpa) 21.66 720.00 15.11 176.80 87.06
FasterVLM (flash attention) 170.30 87.02
@@
FasterVLM (sdpa) 11.51 360.00 14.80 119.87 86.50
FasterVLM (flash attention) 116.79 86.46
@@
FasterVLM (sdpa) 5.71 145.00 14.70 87.36 81.17
FasterVLM (flash attention) 85.26 81.12
@@
FasterVLM (sdpa) 3.80 72.50 14.70 78.18 74.77
FasterVLM (flash attention) 77.66 74.72

### v2

@@ 9 Efficiency Analysis with FlashAttention
VisPruner (sdpa) 6.08 216.00 14.51 101.95 85.92
VisPruner (flash attention) 101.35 85.87
@@
VisPruner (sdpa) 4.16 144.00 14.46 93.57 86.20
VisPruner (flash attention) 92.26 86.16
@@
VisPruner (sdpa) 2.26 72.00 14.44 85.06 83.46
VisPruner (flash attention) 84.03 83.42
@@
VisPruner (sdpa) 1.13 29.00 14.54 79.11 75.85
VisPruner (flash attention) 77.44 75.82
@@
VisPruner (sdpa) 0.76 14.50 14.54 78.09 67.24
VisPruner (flash attention) 77.15 67.22
Table 6: Efficiency comparison between FastV and VisPruner under LLaVA-1.5-7B.
@@
VisPruner (sdpa) 11.50 337.50 26.58 138.36 86.73
VisPruner (flash attention) 137.95 86.72
@@
VisPruner (sdpa) 7.76 225.00 26.31 124.13 86.05
VisPruner (flash attention) 123.33 86.04
@@
VisPruner (sdpa) 4.05 112.50 26.29 104.81 83.10
VisPruner (flash attention) 103.57 83.09
@@
VisPruner (sdpa) 1.86 45.31 26.27 94.80 74.71
VisPruner (flash attention) 94.04 74.66
@@
VisPruner (sdpa) 1.12 22.66 26.43 94.48 65.90
VisPruner (flash attention) 93.73 65.79
Table 7: Efficiency comparison between FastV and VisPruner under LLaVA-1.5-13B.
@@
VisPruner (sdpa) 32.35 1080.00 16.33 246.38 86.69
VisPruner (flash attention) 232.07 86.66
@@
VisPruner (sdpa) 21.66 720.00 15.11 176.80 87.06
VisPruner (flash attention) 170.30 87.02
@@
VisPruner (sdpa) 11.51 360.00 14.80 119.87 86.50
VisPruner (flash attention) 116.79 86.46
@@
VisPruner (sdpa) 5.71 145.00 14.70 87.36 81.17
VisPruner (flash attention) 85.26 81.12
@@
VisPruner (sdpa) 3.80 72.50 14.70 78.18 74.77
VisPruner (flash attention) 77.66 74.72
Table 8: Efficiency comparison between FastV and VisPruner under LLaVA-NeXT-7B.